Suppose the market for watermelons is perfectly competitive and that there are 100 identical firms currently in the market. Each firm as a short run total cost curve of STC=2Q^2+150, with $150 of the fixed costs sunk. calculate the shutdown price for a typical firm.
Answer #1

Answer

A firm will shut down if it fails to recover its total Variable cost i.e. TR < TVC.

TR = Total Revenue == P*Q and TVC = Total Variable cost*Q , P = Price, AVC = Average variable cost.

Here STC = 2Q2 + 150. As 150 is a fixed cost, thus remaining is variable and hence TVC = 2Q2

=> AVC = TVC/Q = 2Q.

In order to maximize profit a perfect competitive firm produces that quantity at which P = MC

where MC = Marginal cost = d(STC)/dQ = 4Q.

Thus P = MC = 4Q and AVC = 2Q. Hence, P > AVC.

Hence there is no P such that P < AVC.

Hence, firm will not shut down at any price.